OpenStudy (anonymous):

how do i write an equations in standard form of (-8,4) and (4,-4)

OpenStudy (gorv):

y-(-4)=(4-(-4))/(-8-4) * (x-4) y-4=8/-12 * (x-4)

OpenStudy (gorv):

@miamiheat17 welcome to openstudy

OpenStudy (gorv):

can u solve further

OpenStudy (anonymous):

THANK YOU

OpenStudy (gorv):

always welcome mate :)